# Visual Structures Helps Visual Reasoning: Addressing the Binding Problem in VLMs

### v2

@@ E.4 Spatial Relationship
Figure 15: An example of a 3D visual search task illustrating the impact of our row-wise structure on the VLM model’s output. The baseline fails to detect the red sphere, returning [False], while applying our method successfully detects it and returns [True].
@@
Figure 17: An example of a 3D counting task and the impact of our proposed method on the VLM model’s output. The baseline method inaccurately counts the number of spheres in the 3D scene, whereas applying our method enables the VLM to accurately count the number of spheres.
@@
Figure 19: An example of a 2D scene description task and the impact of our method on the VLM’s output. The VLM fails to describe the purple star shape in the scene, whereas our method correctly describes all the objects.
Figure 20: An example of a 3D scene description task and the improvement gained by applying our method. The VLM incorrectly adds an extra object, counting three green toruses instead of two, which aren’t part of the scene. In contrast, our proposed method accurately describes only the objects present.
@@
Figure 22: An example of a 3D spatial relationship task and the impact of applying our method on the VLM’s output. As shown in the figure, adding grid lines improves the VLM’s ability to select the correct choice.
Figure 23: An example from a natural dataset for the spatial relationship task and the impact of applying our method on the VLM’s output. The question asks about the spatial relationship between the jar and glasses from the camera’s perspective. As shown in the figure, applying our method helps the VLM answer the question correctly.
@@ Appendix F Attention visualization
Figure 24 shows the results of our method, where we instruct the model to "Scan sequentially based on horizontal lines exist in the image," encouraging it to analyze each row independently. The extracted attention maps for each predicted numerical value confirm this approach, demonstrating that the model attends more strongly to the corresponding spatial regions. For comparison, Figure 25 presents the results from the baseline model, which was given the same instruction. In this case, the attention is less focused, and the predicted numbers do not align clearly with their respective regions. This shows that adding the horizontal line guidance enhanced the model’s ability to focus on each region independently, resulting in improved performance and greater precision.
By applying our method, the model is guided to focus sequentially on smaller, more localized regions within the image. This enables it to more accurately identify and attend to objects in specific areas. The localized approach allows for a more detailed and focused analysis, leading to better alignment between the predicted numerical values and the corresponding spatial regions. In contrast, the baseline method attempts to attend to the entire image in a single step, without such localized guidance. This broader attention scope reduces the model’s ability to focus on specific regions, causing difficulty with precise object identification and reducing accuracy. The model fails to effectively resolve finer details and spatial relationships within the image. To further illustrate this, Figure 26 shows the baseline’s predicted numbers, which attempt to attend to all circles simultaneously. Due to the high number of circles, the model fails to answer the question correctly. Conversely, Figure 24 demonstrates that by incorporating explicit sequential reasoning, our method allows the model to count objects in each region more accurately and aggregate the results more effectively.

### v4

@@ D.4 Spatial Relationship
Figure 16: An example of a 3D visual search task illustrating the impact of our row-wise structure on the VLM model’s output. The baseline fails to detect the red sphere, returning [False], while applying VISER successfully detects it and returns [True].
@@
Figure 18: An example of a 3D counting task and the impact of our proposed method on the VLM model’s output. The baseline method inaccurately counts the number of spheres in the 3D scene, whereas applying VISER enables the VLM to accurately count the number of spheres.
@@
Figure 20: An example of a 2D scene description task and the impact of VISER on the VLM’s output. The VLM fails to describe the purple star shape in the scene, whereas VISER correctly describes all the objects.
Figure 21: An example of a 3D scene description task and the improvement gained by applying VISER. The VLM incorrectly adds an extra object, counting three green toruses instead of two, which aren’t part of the scene. In contrast, our proposed method accurately describes only the objects present.
@@
Figure 23: An example of a 3D spatial relationship task and the impact of applying VISER on the VLM’s output. As shown in the figure, adding grid lines improves the VLM’s ability to select the correct choice.
Figure 24: An example from a natural dataset for the spatial relationship task and the impact of applying VISER on the VLM’s output. The question asks about the spatial relationship between the jar and glasses from the camera’s perspective. As shown in the figure, applying VISER helps the VLM answer the question correctly.
Appendix E Failure case
VISER adds lines in a fixed configuration, independent of image content. While this design ensures consistency and simplicity, it can introduce limitations. For instance, added lines may intersect with important objects, potentially introducing ambiguity about those objects. Alternatively, objects may lie in regions where the added lines provide little to no benefit in decomposing the task. To address this, we conduct two additional controlled experiments on the 2D counting task, targeting conditions where our intervention may help or hurt.
E.1 Object-Line Interaction Analysis
We synthetically vary the level of object-line interference in scenes with circular objects and group data into bins (0.0–0.2: low interference; 0.8–1.0: high interference). As illustrated in Table 10, Our method improves performance significantly in low-interference cases. However, in high-interference scenarios (e.g., dense line-object overlaps), the benefit diminishes. This highlights a key failure mode: when scaffolding introduce ambiguity in object detection. For example, overlapping lines can cause objects to be either double-counted (across both lines) or missed entirely (not clearly belonging to either region), leading to reduced model accuracy. Interestingly, GPT-4o’s baseline accuracy drops with increased density, while Qwen2.5-VL shows the opposite trend. An example of this synthetic dataset is illustrated in Figure 25a.
Table 10: Accuracy of GPT-4o and Qwen2.5-VL on the counting task across different levels of average line-object intersection (ranging from 0.0–0.2 to 0.8–1.0).
Model Method 0.0-0.2 0.2-0.4 0.4-0.6 0.6-0.8 0.8-1.0
GPT-4o Baseline 9.50 12.00 7.00 3.50 3.50
VISER 26.00 26.00 22.00 14.00 13.00
Qwen2.5-VL Baseline 3.00 4.00 10.50 13.50 18.00
VISER 37.00 34.50 17.50 5.00 9.00
E.2 Object Distribution (Entropy) Analysis
To probe another failure case, we conducted a controlled analysis targeting such potential failure modes. Specifically, we evaluated performance with respect to object distributions, which quantifies how widely objects are spread across rows. Low entropy indicates objects are concentrated in one row, while high entropy corresponds to a uniform spread. As illustrated in Table 11, VISER is more effective when objects are distributed across the image (high entropy), where scaffolding provides meaningful visual separation. However, when most objects are located in a single region (low entropy), scaffolds are less helpful, sometimes even harmful. This exposes a second failure case: ineffective intervention when object positioning limits scaffold utility. An example of this synthetic dataset is illustrated in Figure 25b.
Table 11: Accuracy of GPT-4o and Qwen2.5-VL on the counting task across different levels of object spatial entropy (ranging from 0.00–0.50 to 1.75–2.00).
Model Method 0.00-0.50 0.50-1.00 1.00-1.25 1.25-1.50 1.50-1.75 1.75-2.00
GPT-4o Baseline 29.00 30.00 27.00 33.00 38.00 37.00
VISER 48.00 54.00 62.00 62.00 69.00 68.00
Qwen2.5-VL Baseline 14.00 16.00 17.00 16.00 7.00 5.00
VISER 1.00 15.00 16.00 26.00 33.00 36.00
Our primary experiments are conducted on randomly generated datasets with naturally high entropy ( 1.9) and low average interference ( 0.174), where our method is most effective.
Figure 25: Visualization of failure cases in the 2D counting task. Subfigure (a) shows examples with varying levels of object-line interference, while Subfigure (b) illustrates different object spatial distributions.
@@ Appendix F Attention visualization
Figure 26 shows the results of VISER, where we instruct the model to "Scan sequentially based on horizontal lines exist in the image," encouraging it to analyze each row independently. The extracted attention maps for each predicted numerical value confirm this approach, demonstrating that the model attends more strongly to the corresponding spatial regions. For comparison, Figure 27 presents the results from the baseline model, which was given the same instruction. In this case, the attention is less focused, and the predicted numbers do not align clearly with their respective regions. This shows that adding the horizontal line guidance enhanced the model’s ability to focus on each region independently, resulting in improved performance and greater precision.
By applying VISER, the model is guided to focus sequentially on smaller, more localized regions within the image. This enables it to more accurately identify and attend to objects in specific areas. The localized approach allows for a more detailed and focused analysis, leading to better alignment between the predicted numerical values and the corresponding spatial regions. In contrast, the baseline method attempts to attend to the entire image in a single step, without such localized guidance. This broader attention scope reduces the model’s ability to focus on specific regions, causing difficulty with precise object identification and reducing accuracy. The model fails to effectively resolve finer details and spatial relationships within the image. To further illustrate this, Figure 28 shows the baseline’s predicted numbers, which attempt to attend to all circles simultaneously. Due to the high number of circles, the model fails to answer the question correctly. Conversely, Figure 26 demonstrates that by incorporating explicit sequential reasoning, VISER allows the model to count objects in each region more accurately and aggregate the results more effectively.
@@
Appendix G Computational cost
To ensure a fair comparison of computational cost, we measured the average number of output tokens generated across different tasks. Specifically, we compared three methods using the GPT-4o model: the Baseline, our proposed method (VISER), and a Chain-of-Thought (CoT) prompting strategy. Table 12 reports the average token count for each method across four task types, each evaluated on 100 samples.
Table 12: Comparison of average generated tokens for VISER, CoT, and the Baseline across different tasks using the GPT-4o model.
Method Visual Search Counting Scene Desc. Spatial Rel.
Baseline 83.88 11.83 61.77 8.43
CoT 90.48 103.12 62.13 12.52
VISER 153.39 40.94 62.67 10.50